OpenStudy (anonymous):

Find f '(x). f(x) = (x2 + 4x)(x + 7)

OpenStudy (anonymous):

I got 3x^2+22x+20 but was wrong.. I guess I'm not sure where I went wrong..

OpenStudy (rational):

x'the term is 28x right ?

OpenStudy (anteater):

Use product rule, first function times derivative of second function plus second function times derivative of first function. Alternately, you could do the multiplication, then simply use the power rule on the resulting terms.

OpenStudy (rational):

(x2 + 4x)(x + 7) = x^3 + 11x^2 + 28x

OpenStudy (anteater):

(x^2 + 4x)(1) + (2x + 4)(x+7) = x^2 + 4x + 2x^2 + 14x + 4x + 28 = 3x^2 + 22x + 28

OpenStudy (anteater):

Using the product rule

OpenStudy (anteater):

Or you could do what rational did and multiply the expressions together first, then find the derivative of x^3 + 11x^2 + 28x, which should give the same result. I think that way is faster for this problem. :)

OpenStudy (anonymous):

@xydic The +20 should be +28

OpenStudy (anonymous):

Thanks! I have like 20 more questions to answer in 20 mins.. eeeks

OpenStudy (anteater):

Keep 'em coming :D